Evaluate det (A) by expanding along the 2nd row.
1
Expert's answer
2020-05-09T15:34:12-0400
"det(A)=\sum_{j=1}^n (-1)^{i+j} a_{ij} M_j^i"

"M_j^i - additional \space minor \space to \space element \space a_{ij}"


"det(A)=\begin{vmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{21} & a_{22} & a_{23} \\
 a_{31} & a_{32} & a_{33}
\end{vmatrix}
=(-1)^{2+1}a_{21}\begin{vmatrix}
 a_{12} & a_{13} \\
 a_{32} & a_{33}
\end{vmatrix}+"


"(-1)^{2+2}a_{22}\begin{vmatrix}
 a_{11} & a_{13} \\
 a_{31} & a_{33}
\end{vmatrix}+(-1)^{2+3}a_{23}\begin{vmatrix}
 a_{11} & a_{12} \\
 a_{31} & a_{32}
\end{vmatrix}="

"-a_{21}(a_{12}a_{33}-a_{13}a_{32})+a_{22}(a_{11}a_{33}-a_{13}a_{31})-a_{23}(a_{11}a_{32}-a_{12}a_{31})="


"-a_{12}a_{21}a_{33}+a_{13}a_{21}a_{32}+a_{11}a_{22}a_{33}-a_{13}a_{22}a_{31}-a_{11}a_{23}a_{32}+a_{12}a_{23}a_{31}"